In January 2014 a massive rockfall in New Zealand reduced the height of Aoraki, or Mt. Cook, from the previously measured 12,316 feet to 12,218 feet. Determine the difference in meters of the height of Mt. Cook before and after rockfall. And determine percent of height lost by Mt.Cook due to rockfall. Thanks in advance
Answer #1

Difference in feet of the height of Mt. Cook before and after rockfall = 12316 feet - 12218 feet = 98 feet

We have relation, 1 m = 3.28084 feet. Conversion factor is 1 m / 3.28084 feet.

  Difference in meters of the height of Mt. Cook before and after rockfall = 98 feet x ( 1 m / 3.28084 feet.) = 29.870 m

% height lost = (Height lost / original height) x 100 = ( 98/ 12316) x 100 = 0.8

% height lost by Mt. cook due to rockfall = 0.8